# Pyramid Vector Quantization for LLMs

### v1

@@ Abstract
Recent works on compression of large language models (LLM) using quantization considered reparameterizing the architecture such that weights are distributed on the sphere. This demonstratively improves the ability to quantize by increasing the mathematical notion of coherence, resulting in fewer weight outliers without affecting the network output. In this work, we aim to further exploit this spherical geometry of the weights when performing quantization by considering Pyramid Vector Quantization (PVQ) for large language models. Arranging points evenly on the sphere is notoriously difficult, especially in high dimensions, and in case approximate solutions exists, representing points explicitly in a codebook is typically not feasible due to its additional memory cost. Instead, PVQ uses a fixed integer lattice on the sphere by projecting points onto the 1-sphere, which allows for efficient encoding and decoding without requiring an explicit codebook in memory. To obtain a practical algorithm, we propose to combine PVQ with scale quantization for which we derive theoretically optimal quantizations, under empirically verified assumptions. Further, we extend pyramid vector quantization to use Hessian information to minimize quantization error under expected feature activations, instead of only relying on weight magnitudes. Experimentally, we achieves state-of-the-art quantization performance with pareto-optimal trade-off between performance and bits per weight and bits per activation, compared to compared methods. On weight-only, we find that we can quantize a Llama-3 70B model to 3.25 bits per weight and retain 98% accuracy on downstream tasks.
@@ 1 Introduction
PVQ has been been very successful in well-known audio (Valin et al., 2012) and video codecs (Daede et al., 2016). We demonstrate that the same algorithm allows practical quantization of large language models, by proposing a group-wise quantization scheme and further extending PVQ to use Hessian information accounting for curvature in the loss. We also propose a scheme to quantize the normalized scales (gains) of each group according to theoretically derived quantiles, which we verified to closely match the empirical weight distributions of pretrained LLMs in practice. Our contribution extends beyond Liguori (2017), who considered PVQ on weights of small neural networks.
@@ 2.1 Quantization
$L(\widehat{{\bm{W}}})=\mathbb{E}_{{\bm{x}}}\left[||{\bm{W}}{\bm{x}}-\widehat{{\bm{W}}}{\bm{x}}||_{2}\right]=\mathbb{E}_{{\bm{x}}}\left[\operatorname{Tr}\left(({\bm{W}}-\widehat{{\bm{W}}}){\bm{H}}({\bm{W}}-\widehat{{\bm{W}}}\right)\right]$ (1)
where the layer-wise Hessian ${\bm{H}}=\mathbb{E}_{{\bm{x}}}\left[{\bm{x}}{\bm{x}}^{T}\right]=\sum_{n=1}^{N}{\bm{x}}_{n}{\bm{x}}_{n}^{T}$ is typically empirically approximated using a calibration dataset $\{{\bm{x}}_{n}\}_{n=1}^{N}$. The objective is optimal in the sense that it minimizes the layer-wise output at each layer, but a crude approximation with respect to the actual training loss. Some recent works have proposed to also use gradients to improve the approximation to the true training objective (van der Ouderaa et al., 2023). Although it would be interesting to combine ideas presented in this paper with gradients, we stick to a layer-wise loss Equation 1 for simplicity and because this yields a faster quantization method that does not require backpropagation.
@@ 2.2 Vector quantization
Instead of individually quantizing weights, as done in scalar quantization, vector quantization aims to simultaneously quantize multiple weights. It can be shown that, even for completely independent Gaussian sources, this typically results in much higher theoretical signal-to-noise ratios, leading recent works to consider vector quantization for LLMs (van Baalen et al., 2024; Liu et al., 2024b). Yet, vector quantization is not widely adopted in practice because of two practical problems. Firstly, naive vector quantization requires constructing an explicit codebook using clustering (such as K-means), quickly becoming infeasibly large for higher number of dimensions. Secondly, quantization requires an expensive search over this explicit codebook, which can not practically be used on-the-fly on activations. This limits VQ to weights-only quantisation. Pyramid vector quantization is a vector quantization method that solves these issues simultaneously by being codebook-free and search-free, unlike standard vector quantization. It uses an implicit codebook that does not have to be instantiated in memory to perform quantization, and can encode and decode vectors without having to search it. As a result, it can be applied on-the-fly and not only on the neural network weights, but also to the activations during inference.
@@ 2.3 Weights on the sphere
Instead of viewing weight vectors in Euclidean $D$-space ${\bm{w}}\in\mathbb{R}^{D}$, they can be interpreted as scaled points ${\bm{w}}=s{\bm{v}}$ on the unit sphere ${\bm{v}}\in\Omega_{D}$, with $\Omega_{D}=\{{\bm{v}}=(v_{1},v_{2},\ldots,v_{D})\in\mathbb{R}^{D}:\|{\bm{v}}\|_{2}=1\}$. We refer to this spherical decomposition as the shape ${\bm{v}}=\frac{{\bm{w}}}{\|{\bm{w}}\|_{2}}$ and the gain $s=\|{\bm{w}}\|_{2}$ of a vector. Recent works that use such a spherical perspective on LLM weights have offered new insights in properties of the training dynamics and guide algorithmic improvements. For example, (Kosson et al., 2023) noted that LLM weights under weight decay or popular deep learning optimizers converge to an equilibrium on the sphere, theoretically predicting the magnitude of the gain after training. Some recent works have shown that only updating shape components can be beneficial in low-rank adaptation Liu et al. (2024a) and training itself (Loshchilov et al., 2024). Yet, weights are not always uniformly distributed making quantization hard and giving rise to outliers. We use recent work on coherence processing that reparameterize weights on the sphere in a way that reduces outliers without functionally changing the output of the network. We then extend existing quantization approaches by considering pyramid vector quantization to construct a spherical quantization code.
@@
Recent works have shown that by considering rotations of the weight matrices ${\bm{W}}\in\mathbb{R}^{R\times C}$, it is possible to improve weight coherence Chee et al. (2024), thereby diminishing the number of outliers without altering the network’s output, ${\bm{W}}={\bm{U}}{\bm{W}}{\bm{V}}$ where ${\bm{U}}\in\mathbb{R}^{R\times R}$ and ${\bm{V}}\in\mathbb{R}^{C\times C}$ are both orthogonal matrices. To perform a forward pass this reparameterization must be undone, $\widehat{W}={\bm{V}}^{T}\widehat{{\bm{W}}}{\bm{U}}^{T}$. To make this efficient, Chee et al. (2024) proposed to use random Hadamard matrices for ${\bm{U}}$ and ${\bm{V}}$.
@@
In most cases, the additional rotation matrices do not result in a memory or compute overhead during inference since rotation matrices can be absorbed into weight matrices: depending on the placement of the rotation matrix in the architecture, it can be left- or right- multiplied with an adjacent weight matrix. This principle was used in QuaRot (Ashkboos et al., 2024), to which we refer for a more detailed description of this for attention and fully-connected layers found in most common LLM architectures including specifics around absorbing rotation matrices in the presence of activation functions and normalization layers.
@@ 2.4 Classic Pyramid Vector Quantization
Figure 1: Illustration of the PVQ integer lattice in $d=3$ dimensions with increasing pulses $k$ from 1 to 6. Points on the pyramid $P_{3,k}$ are projected onto the sphere $S_{3,k}$.
To quantize points on the sphere, we would like to construct spherical code, a finite subset on the unit sphere $S\subset\Omega_{D}$. Packing a set of points on the surface of a sphere such that their distance is maximised is a notoriously hard problem in mathematics, famously dating back to the Dutch botanist Tammes (1930). Just like sphere packing, optimal spherical codes are not generally known, with the exception of some specific dimensions, similar to the E8 packing in 8 dimensions exploited in (Tseng et al., 2024). Even though good but sub-optimal spherical codes exist (Conway & Sloane, 2013), there is not always an efficient method to enumerate the packing without requiring an explicit codebook, making use for quantization costly. PVQ (Fischer, 1986) provides a solution to both of these issues simultaneously, allowing good spherical codes to be constructed in arbitrary dimension that can be efficiently encoded and decoded without having to maintain an explicit codebook in memory. This is achieved by projecting an integer lattice on the L1 ball $P_{D,K}$ onto the hypersphere $S_{D,K}$. We now provide a brief overview of the classic PVQ algorithm.
@@
Formally, the integer lattice of PVQ on the $D$-dimensional hyper-sphere is obtained by starting from a set of points on the L1 ball of radius $K$, and projecting the set of points onto the sphere. We denote the set of integer points on the L1 ball as $P_{D,K}$,
$P_{D,K}=\left\{{\bm{p}}\in\mathbb{Z}^{D}:||{\bm{p}}||_{1}=\sum_{d=1}^{D}|{\bm{p}}_{d}|=K\right\}$ (2)
To obtain our spherical code $S_{D,K}$, we can project the points $P_{D,K}$ onto the sphere, by normalizing,
$S_{D,K}=\left\{\frac{{\bm{p}}}{||{\bm{p}}||_{2}}:{\bm{p}}\in P_{D,K}\right\}\,.$ (3)
The number of codes $N(D,K)=|S_{D,K}|=|P_{D,K}|$ can be written as
@@ 2.5 Subroutines of classic PVQ
Quantizing the shape
To quantizing a vector ${\bm{w}}\in\mathbb{R}^{D}$, we map it to the closest point on the pyramid $P_{D,K}$ using an iterative procedure that projects the vector ${\bm{w}}_{0}={\bm{w}}$ onto the L1 ball of radius $K$, and round it to the closest integer,
@@
Figure 2: Illustration of points on pyramid $P_{2,7}$, their projections onto the sphere $S_{2,7}$ and codes in $C_{2,7}$.
@@
We can encode points on the pyramid ${\bm{p}}\in P_{D,K}$ to integer codes $c\in[1,\ldots,N(D,K)-1]$ using an algorithm, which avoids having to build an explicit table in memory to represent the codebook. The original PVQ paper (Fischer, 1986) describes an encoding scheme, of which we provide pseudocode Algorithm 2 in Section B.2, for the purpose of being self-contained. The algorithm provides a bijective mapping from the points of the pyramid $P_{D,K}$ to integer indices $C_{D,K}$, providing a compact representation that allow vectors to be efficiently stored in as few bits as possible. In Figure 2, we provide an example illustration of points on $P_{2,7}$ and indices $C_{2,7}=[0,27]$. The table can be precomputed a table of $N(d,k)$ for $0\leq d\leq D$ and $0\leq k\leq K$ and reused to avoid recomputing the same quantity.
@@
We decode the integer codes $c\in C$ of PVQ back to the associated vector ${\bm{p}}\in P_{D,K}$ by performing a decoding algorithm which performs the inverse operation of the encoding algorithm above. In Section B.3, we provide a corrected version of the decoding algorithm described in the original PVQ paper (Fischer, 1986). The original paper contains an error and misses a line after setting $x_{i}\leftarrow-j$, resulting in wrong decodings when ${\bm{p}}$ contains negative values, except for when the last value is negative (which happens to be the case for the example given in the original paper).
@@ 3 PVQ for LLM compression
3.1 Why use PVQ for LLM quantization?
PVQ offers several practical advantages over competing methods. Firstly, PVQ is a vector quantization method which means it can achieve higher signal-to-noise ratios than scalar quantization methods that quantize weights independently. Secondly, PVQ is codebook-free, which means that it does not require an explicit codebook to be constructed in memory. This makes the approach more memory efficient, but more importantly, the implicit codebook size can reach far beyond the memory that would have been required with an explicit codebook. To illustrate, explicitly storing a codebook that quantizes 16 bit precision vectors of a groupsize of 128 to 4 bit per weight would require approximately $2.7\cdot 10^{154}$ bytes, exceeding the estimated information capacity of the observable universe. With PVQ, we can use implicit codebooks of this size because encoding and decoding are done by an algorithm, not a table lookup. Thirdly, PVQ is search-free, which means that vectors can be encoded without having to search over a codebook. This is significant apart from the computational benefits, because it allows for on-the-fly quantization of activations and opens the door to quantization at train time. Lastly, the desired bits per weight after quantization can also be fractions (e.g. $b_{\text{shape}}=3.5$) and are not limited to integers. The ability to choose the groupsize, and bits for the shape make PVQ highly flexible and can be chosen such that the most optimal trade-off between compression and performance is achieved. We find that PVQ outperforms competitive quantization methods in terms of weight-only and activation quantization.

### v2

@@ Abstract
Recent works on compression of large language models (LLM) using quantization considered reparameterizing the architecture such that weights are distributed on the sphere. This demonstratively improves the ability to quantize by increasing the mathematical notion of coherence, resulting in fewer weight outliers without affecting the network output. In this work, we aim to further exploit this spherical geometry of the weights when performing quantization by considering Pyramid Vector Quantization (PVQ) for large language models. Arranging points evenly on the sphere is notoriously difficult, especially in high dimensions, and in case approximate solutions exists, representing points explicitly in a codebook is typically not feasible due to its additional memory cost. Instead, PVQ uses a fixed integer lattice on the sphere by projecting points onto the 1-sphere, which allows for efficient encoding and decoding without requiring an explicit codebook in memory. To obtain a practical algorithm, we propose to combine PVQ with scale quantization for which we derive theoretically optimal quantizations, under empirically verified assumptions. Further, we extend pyramid vector quantization to use Hessian information to minimize quantization error under expected feature activations, instead of only relying on weight magnitudes. Experimentally, we achieve state-of-the-art quantization performance with pareto-optimal trade-off between performance and bits per weight and bits per activation, compared to competitive methods. On weight-only, we find that we can quantize a Llama-3 70B model to 3.25 bits per weight and retain 98% accuracy on downstream tasks.
@@ 1 Introduction
PVQ has been very successful in well-known audio (Valin et al., 2012) and video codecs (Daede et al., 2016). We demonstrate that the same algorithm allows practical quantization of large language models, by proposing a group-wise quantization scheme and further extending PVQ to use Hessian information accounting for curvature in the loss. We also propose a scheme to quantize the normalized scales (amplitudes) of each group according to theoretically derived quantiles, which we verified to closely match the empirical weight distributions of pretrained LLMs in practice. Our contribution extends beyond Liguori (2017), which considered PVQ on weights of small neural networks.
@@ 2.1 Quantization
$L(\widehat{{\bm{W}}})=\mathbb{E}_{{\bm{x}}}\left[||{\bm{W}}{\bm{x}}-\widehat{{\bm{W}}}{\bm{x}}||_{2}\right]=\operatorname{Tr}\left(({\bm{W}}-\widehat{{\bm{W}}}){\bm{H}}({\bm{W}}-\widehat{{\bm{W}}})\right)$ (1)
where the layer-wise Hessian ${\bm{H}}=\mathbb{E}_{{\bm{x}}}\left[{\bm{x}}_{n}{\bm{x}}_{n}^{T}\right]\approx\frac{1}{N}\sum_{n=1}^{N}{\bm{x}}_{n}{\bm{x}}_{n}^{T}$ is empirically estimated using a calibration dataset $\{{\bm{x}}_{n}\}_{n=1}^{N}$. The objective is optimal in the sense that it minimizes the layer-wise output at each layer, but a crude approximation with respect to the actual training loss. Some recent works have proposed to also use gradients to improve the approximation to the true training objective (van der Ouderaa et al., 2023). Although it would be interesting to combine ideas presented in this paper with gradients, we stick to a layer-wise loss Equation 1 for simplicity and because this yields a faster quantization method that does not require backpropagation.
@@ 2.2 Vector quantization
Instead of individually quantizing weights, as done in scalar quantization, vector quantization aims to simultaneously quantize multiple weights. It can be shown that, even for completely independent Gaussian sources, this typically results in much higher theoretical signal-to-noise ratios, leading recent works to consider vector quantization for LLMs (van Baalen et al., 2024; Liu et al., 2024b; Egiazarian et al., 2024; Tseng et al., 2024a; b). Yet, vector quantization is not widely adopted in practice because of two practical problems. Firstly, naive vector quantization requires constructing an explicit codebook using clustering (such as K-means), quickly becoming infeasibly large for higher number of dimensions. Secondly, quantization requires an expensive search over this explicit codebook, which can not practically be used on-the-fly on activations. Although application to LLMs is limited, it is common in vector quantization (Gray & Neuhoff, 1998) to use an implicit codebook which does not have to be explicitly instantiated in memory. PVQ is such a vector quantization and comes with the additional benefit of being search-free, allowing encoding and decoding of vectors without having to perform an explicit and exhaustive lookup that at least linearly scales in algorithmic complexity with the size of the codebook. As a result, it can be applied on-the-fly and not only on the neural network weights, but also to the activations during inference.
@@ 2.3 Weights on the sphere
Instead of viewing weight vectors in Euclidean $D$-space ${\bm{w}}\in\mathbb{R}^{D}$, they can be interpreted as scaled points ${\bm{w}}=s{\bm{v}}$ on the unit sphere ${\bm{v}}\in\Omega_{D}$, with $\Omega_{D}=\{{\bm{v}}=(v_{1},v_{2},\ldots,v_{D})\in\mathbb{R}^{D}:\|{\bm{v}}\|_{2}=1\}$. We refer to this spherical decomposition as the direction ${\bm{v}}=\frac{{\bm{w}}}{\|{\bm{w}}\|_{2}}$ and the amplitude $s=\|{\bm{w}}\|_{2}$ of a vector. Recent works that use such a spherical perspective on LLM weights have offered new insights in properties of the training dynamics and guide algorithmic improvements. For example, (Kosson et al., 2023) noted that LLM weights under weight decay or popular deep learning optimizers converge to an equilibrium on the sphere, theoretically predicting the magnitude of the amplitude after training. Recent works have shown that only updating direction components can be beneficial in low-rank adaptation Liu et al. (2024a) and training itself (Loshchilov et al., 2024). Yet, weights are not always uniformly distributed making quantization hard and giving rise to outliers. We use coherence processing (Chee et al., 2024) to reparameterize weights on the sphere in a way that reduces outliers without functionally changing the output of the network. Further, in high dimensions the information in the amplitude is negligible compared to the information in the direction (Kipnis & Reeves, 2021). The observation that LLM weights are in practice uniformly distributed across the sphere is the primary motivation behind exploring pyramid vector quantization, which allows us to construct an efficient spherical quantization code.
@@
Recent works have shown that rotating weights (Chee et al., 2024; Ashkboos et al., 2024) can improve weight coherence and reduce the number of outliers without altering the network’s output, $\widetilde{{\bm{W}}}={\bm{U}}{\bm{W}}{\bm{V}}$ where ${\bm{U}}\in\mathbb{R}^{R\times R}$ and ${\bm{V}}\in\mathbb{R}^{C\times C}$ are both orthogonal matrices. Since the transpose of orthogonal matrices equals their inverse, the reparameterisation can simply be undone in the forward pass by left- and right multiplying with matrix transposes, ${\bm{W}}={\bm{U}}^{T}\widetilde{{\bm{W}}}{\bm{V}}^{T}$. We follow Chee et al. (2024); Ashkboos et al. (2024) and use random Hadamard matrices for ${\bm{U}}$ and ${\bm{V}}$, which can be implemented very efficiently.
@@
In most cases, the additional rotation matrices do not result in a memory or compute overhead during inference since rotation matrices can be absorbed into weight matrices: depending on the placement of the rotation matrix in the architecture, it can be left- or right- multiplied with an adjacent weight matrix. This principle was used in QuaRot (Ashkboos et al., 2024), which describes how rotation matrices can be efficiently absorbed into attention and fully-connected layers in commonly used LLM architectures.
@@ 2.4 Classic Pyramid Vector Quantization
Figure 1: Illustration of the PVQ integer lattice in $d=3$ dimensions with increasing pulses $k$ from 1 to 6. Points on the pyramid ${\mathcal{P}}_{3,k}$ are projected onto the sphere ${\mathcal{S}}_{3,k}$.
To quantize points on the sphere, we would like to construct spherical code, a finite subset on the unit sphere ${\mathcal{S}}\subset\Omega_{D}$. Packing a set of points on the surface of a sphere such that their distance is maximized is a notoriously hard problem in mathematics, famously dating back to the Dutch botanist Tammes (1930). Just like sphere packing, optimal spherical codes are not generally known, with the exception of some specific dimensions, similar to the E8 packing in 8 dimensions exploited in (Tseng et al., 2024a). Even though good but sub-optimal spherical codes exist (Conway & Sloane, 2013), there is not always an efficient method to enumerate the packing without requiring an explicit codebook, resulting in costly quantization. PVQ (Fischer, 1986) provides a solution to both of these issues simultaneously, allowing good spherical codes to be constructed in arbitrary dimension that can be efficiently encoded and decoded without having to maintain an explicit codebook in memory. This is achieved by projecting an integer lattice on the $l_{1}$ ball ${\mathcal{P}}_{D,K}$ onto the hypersphere ${\mathcal{S}}_{D,K}$. We now formalize these concepts by providing an overview of classic PVQ algorithm.
@@
Formally, the integer lattice of PVQ on the $D$-dimensional hyper-sphere is obtained by starting from a set of points on the $l_{1}$ ball of radius $K$, and projecting the set of points onto the sphere. We denote the set of integer points on the $l_{1}$ ball as ${\mathcal{P}}_{D,K}$,
${\mathcal{P}}_{D,K}=\left\{{\bm{p}}\in\mathbb{Z}^{D}:||{\bm{p}}||_{1}=\sum_{d=1}^{D}|{\bm{p}}_{d}|=K\right\}$ (2)
To obtain our spherical code ${\mathcal{S}}_{D,K}$, we can project the points ${\mathcal{P}}_{D,K}$ onto the sphere, by normalizing,
${\mathcal{S}}_{D,K}=\left\{\frac{{\bm{p}}}{||{\bm{p}}||_{2}}:{\bm{p}}\in{\mathcal{P}}_{D,K}\right\}\,.$ (3)
The number of codes $N(D,K)=|{\mathcal{S}}_{D,K}|=|{\mathcal{P}}_{D,K}|$ can be written as
@@ 2.5 Subroutines of classic PVQ
Quantizing the direction
To quantizing a vector ${\bm{w}}\in\mathbb{R}^{D}$, we map it to the closest point on the pyramid ${\mathcal{P}}_{D,K}$ using an iterative procedure that projects the vector ${\bm{w}}_{0}={\bm{w}}$ onto the $l_{1}$ ball of radius $K$, and round it to the closest integer,
@@
Figure 2: Illustration of points on pyramid ${\mathcal{P}}_{2,7}$, their projections onto the sphere ${\mathcal{S}}_{2,7}$ and codes in ${\mathcal{C}}_{2,7}$.
@@
We can encode points on the pyramid ${\bm{p}}\in{\mathcal{P}}_{D,K}$ to integer codes $c\in[1,\ldots,N(D,K)-1]$ using an efficient algorithm, which avoids having to build an explicit table in memory to represent the codebook. The original PVQ paper (Fischer, 1986) describes an encoding scheme, of which we provide pseudocode Algorithm 2 in Section B.2, for the purpose of being self-contained. The algorithm provides a bijective mapping from the points of the pyramid ${\mathcal{P}}_{D,K}$ to the set of integer indices ${\mathcal{C}}_{D,K}$, providing a compact representation that allow vectors to be efficiently stored in as few bits as possible. In Figure 2, we provide an example illustration of points on ${\mathcal{P}}_{2,7}$ and indices ${\mathcal{C}}_{2,7}=[0,27]$. The table of $N(d,k)$ can be precomputed for $0\leq d\leq D$ and $0\leq k\leq K$ and reused to avoid recomputing the same quantity.
@@
We decode the integer codes $c\in{\mathcal{C}}$ of PVQ to their associated vectors ${\bm{p}}\in{\mathcal{P}}_{D,K}$ through a decoding algorithm which performs the inverse operation of the encoding algorithm above. In Section B.3, we provide a corrected version of the decoding algorithm described in the original PVQ paper (Fischer, 1986). The original paper contains an error and misses a line after setting $x_{i}\leftarrow-j$, resulting in wrong decodings when ${\bm{p}}$ contains negative values, except for when the last value is negative (which happens to be the case for the example given in the original paper).
@@ 3 PVQ for LLM compression
3.1 Practical advantages of PVQ
PVQ offers several practical advantages over competing methods. Firstly, PVQ is a vector quantization method which means it can achieve higher signal-to-noise ratios than scalar quantization methods that quantize weights independently. Secondly, PVQ uses an implicit codebook, which means that it does not require an explicit codebook to be constructed in memory. This makes the approach more memory efficient, but more importantly, the implicit codebook size can reach far beyond the memory that would have been required with an explicit codebook. To illustrate, explicitly storing a codebook that quantizes 16 bit precision vectors of a groupsize of 128 to 4 bit per weight would require approximately $2.7\cdot 10^{154}$ bytes, exceeding the estimated information capacity of the observable universe. With PVQ, we can use implicit codebooks of this size because encoding and decoding are done by an efficient algorithm, not a table lookup. Thirdly, PVQ is search-free, which means that vectors can be encoded without having to search over a codebook. This is significant apart from the computational benefits, because it allows for on-the-fly quantization of activations and opens the door to quantization at train time. Lastly, the desired bits per weight after quantization can also be fractions (e.g. $b_{\text{direction}}=3.5$) and are not limited to integers. The ability to choose the groupsize, and bits for the direction make PVQ highly flexible and can be chosen such that the most optimal trade-off between compression and performance is achieved. We find that PVQ outperforms competitive quantization methods in terms of weight-only and activation quantization.